# Adjusted Connections I: Differential Cocycles for Principal Groupoid Bundles with Connection

### v1

@@ 4.3 Construction of adjustment groupoids
There is now a simple condition which ensures that the pre-adjustment is indeed an adjustment; recall Definition 3.10 for the following.
Theorem 4.14.
Given the pre-adjustment groupoid $\mathscr{A}^{\mathsf{W}}_{\mathcal{H}}$ of a Lie groupoid $\mathscr{G}$ and an adjustment $\nabla$ on its Lie algebroid $\mathfrak{g}$, then $\mathscr{A}^{\mathsf{W}}_{\mathcal{H}}$ is an adjustment if the Lie algebroid connection on $\mathfrak{g}$ is induced by the Cartan connection $\mathcal{H}$ on $\mathscr{G}$ by Proposition 2.12.
In this case we write $\mathscr{A}_{\mathcal{H}}^{\mathsf{W}}$ instead of $\mathscr{A}_{\mathcal{H}}^{\mathsf{W}}$.
Proof.
Recall from (4.8) that we have to show that the presheaves
$\mathcal{N}\mapsto\underline{\mathsf{hom}}^{\rm red}_{\nabla}(\mathcal{N},\mathrm{T}[1]\mathfrak{g}[1])~{}~{}~{}\mbox{and}~{}~{}~{}\mathcal{N}\mapsto\mathsf{Lie}\left(\mathsf{dgGrpd}\left(\mathcal{N}\rightrightarrows\mathcal{N},\mathscr{A}^{\mathsf{W}}_{\mathcal{H}}\right)\right)$ (4.37)
[…]
  • a topological map
    $|f_{\flat}|:|\mathcal{X}|\times|\mathcal{N}|\rightarrow M$ (4.38)
  • and two maps of modules over $|f|^{*}C^{\infty}(M)\subset C^{\infty}(\mathcal{X})_{0}\otimes C^{\infty}(\mathcal{N})_{0}=C^{\infty}(|\mathcal{X}|\times|\mathcal{N}|)$,
    $f^{\sharp}_{\flat 0}:$ $|f|^{*}\Gamma(\mathfrak{g}[1])^{*}$ $\rightarrow C^{\infty}(\mathcal{X})_{0}\otimes C^{\infty}(\mathcal{N})_{1}~{},$ (4.39)
    $f^{\sharp}_{\flat 1}:$ $|f|^{*}\Gamma(\mathfrak{g}[1])^{*}$ $\rightarrow C^{\infty}(\mathcal{X})_{1}\otimes C^{\infty}(\mathcal{N})_{0}~{},$
    subject to the conditions (4.41) in the local coordinates $(m^{a},\bar{m}^{a},\xi^{\alpha},\bar{\xi}^{\alpha})$.
The conditions arise to ensure that $f^{\sharp}\circ\mathsf{p}_{\rm curv}^{\sharp}$ only takes values in $C^{\infty}(\mathcal{X})_{0}\otimes C^{\infty}(\mathcal{N})$. The lift $f$ can be explicitly calculated by imposing that $f$ is a dg-map. In the local coordinates $(m^{a},\bar{m}^{a},\xi^{\alpha},\bar{\xi}^{\alpha})$ and the structure functions introduced in (2.51), $f$ is given by $|f|=|f_{\flat}|$ and
$\begin{split}\tilde{f}^{\sharp}(m^{a})&=\tilde{f}^{\sharp}_{\flat}(m^{a})\\ &=|f_{\flat}|^{*}(m^{a})~{},\\ \tilde{f}^{\sharp}(\xi^{\alpha})&=\tilde{f}^{\sharp}_{\flat}(\xi^{\alpha})\\ &=\tilde{f}^{\sharp}_{\flat 0}(\xi^{\alpha})+\tilde{f}^{\sharp}_{\flat 1}(\xi^{\alpha})~{},\\ \tilde{f}^{\sharp}(\bar{m}^{a})&=(\mathrm{d}_{\mathcal{X}}+\mathrm{d}_{\mathcal{N}})\tilde{f}^{\sharp}_{\flat}(m^{a})-\tilde{f}^{\sharp}_{\flat}(\mathrm{d}_{\mathfrak{g}[1]}(m^{a}))\\ &=(\mathrm{d}_{\mathcal{X}}+\mathrm{d}_{\mathcal{N}})\tilde{f}^{\sharp}_{\flat}(m^{a})-|f_{\flat}|^{*}\mathtt{r}^{a}_{\alpha}\left(\tilde{f}^{\sharp}_{\flat 0}(\xi^{\alpha})+\tilde{f}^{\sharp}_{\flat 1}(\xi^{\alpha})\right)~{},\\ \tilde{f}^{\sharp}(\bar{\xi}^{\alpha})&=(\mathrm{d}_{\mathcal{X}}+\mathrm{d}_{\mathcal{N}})\tilde{f}^{\sharp}_{\flat}(\xi^{\alpha})-\tilde{f}^{\sharp}_{\flat}(\mathrm{d}_{\mathfrak{g}[1]}(\xi^{\alpha}))~{},\end{split}$ (4.40)
where $\tilde{f}_{i}^{\sharp}=f_{i}^{\sharp}\circ|f|^{*}$. For $f^{\sharp}\circ\mathsf{p}_{\rm curv}^{\sharp}$ to only take values in $C^{\infty}(\mathcal{X})_{0}\otimes C^{\infty}(\mathcal{N})$, we need to impose the following relations:
$\begin{gathered}\mathsf{d}_{\mathcal{X}}|f|^{*}h=\left(\tilde{f}^{\sharp}_{1}\circ\rho^{*}\right)(\mathrm{d}h)~{},\\ \mathsf{d}_{\mathcal{X}}\tilde{f}_{0}^{\sharp}\left(\xi^{\alpha}\right)+|f|^{*}\left(\mathtt{f}_{\beta\gamma}^{\alpha}\right)\tilde{f}^{\sharp}_{1}\left(\xi^{\beta}\right)\tilde{f}^{\sharp}_{0}\left(\xi^{\gamma}\right)+|f|^{*}\left(\omega_{a\beta}^{\alpha}\right)\tilde{f}^{\sharp}\left(\bar{m}^{a}\right)\tilde{f}^{\sharp}_{1}\left(\xi^{\beta}\right)+\mathsf{d}_{\mathcal{N}}\tilde{f}^{\sharp}_{1}\left(\xi^{\alpha}\right)=0~{},\\ \mathsf{d}_{\mathcal{X}}\tilde{f}^{\sharp}_{1}\left(\xi^{\alpha}\right)=-\tfrac{1}{2}|f|^{*}\left(\mathsf{f}_{\beta\gamma}^{\alpha}\right)\tilde{f}^{\sharp}_{1}\left(\xi^{\beta}\right)\tilde{f}^{\sharp}_{1}\left(\xi^{\gamma}\right)\end{gathered}$ (4.41)
for all $h\in C^{\infty}(M)$.
that factor through another groupoid:
where the topological part $|\mathsf{p}_{0}|$ of $\mathsf{p}_{0}$ is identity and $\mathsf{p}_{0}^{\sharp}$ is the inclusion of degree 00 elements of the local ring of functions to the local rings of functions and $\check{F}$ is some morphism of graded manifolds. In the following, we will show that such functors are given by
  • a topological map
    $|F|:|\mathcal{X}|\times|\mathcal{N}|\to M$ (4.45)
  • and two maps of modules over $|F|^{*}C^{\infty}(M)$
    $(\bar{F}_{0})^{\sharp}:$ $|F|^{*}\Gamma(\mathfrak{g}[1])^{*}$ $\to C^{\infty}(\mathcal{X})_{0}\otimes C^{\infty}(\mathcal{N})_{1}~{},$ (4.46)
    $(\bar{F}_{1})^{\sharp}:$ $|F|^{*}\Gamma(\mathfrak{g}[1])^{*}$ $\to C^{\infty}(\mathcal{X})_{1}\otimes C^{\infty}(\mathcal{N})_{0}~{},$
A general functor $F$ is given by a dg-map between objects $F_{0}:\mathcal{X}\times\Theta\times\mathcal{N}\to\mathrm{T}[1]\mathfrak{g}[1]$ and a dg-map between morphisms $F_{1}:\mathcal{X}\times\Theta\times\Theta\times\mathcal{N}\to\mathrm{T}[1]\mathfrak{g}[1]\times_{\mathrm{T}[1]M}\mathrm{T}[1]\mathsf{G}$ satisfying functoriality conditions. The dg-map $F_{0}:\mathcal{X}\times\Theta\times\mathcal{N}\rightarrow\mathrm{T}[1]\mathfrak{g}[1]$ is determined by uniquely lifting a graded map $F_{\flat 0}:\mathcal{X}\times\Theta\times\mathcal{N}\rightarrow\mathfrak{g}[1]$ which is given by a topological map $|F_{\flat 0}|=|F|$ as well as maps
$\left(F_{\flat 0}^{\sharp}\right)_{0}:$ $|F|^{*}C^{\infty}(M)$ $\to C^{\infty}(\mathcal{X})_{0}\otimes C^{\infty}(\mathcal{N})_{0}~{}\oplus~{}\theta C^{\infty}(\mathcal{X})_{1}\otimes C^{\infty}(\mathcal{N})_{0}~{},$ (4.47)
$\left(F_{\flat 0}^{\sharp}\right)_{1}:$ $|F|^{*}\Gamma(\mathfrak{g}[1])^{*}$ $\to C^{\infty}(\mathcal{X})_{0}\otimes C^{\infty}(\mathcal{N})_{1}~{}\oplus~{}\theta C^{\infty}(\mathcal{X})_{1}\otimes C^{\infty}(\mathcal{N})_{1}~{},$
where $(F_{\flat 0}^{\sharp})_{0}$ is a ring homeomorphism, $(F_{\flat 0}^{\sharp})_{1}$ is a morphism of modules over $|F|^{*}C^{\infty}(M)$ and $\theta$ is the generator of $C^{\infty}(\Theta)$ with $\mathrm{d}_{\Theta}\theta=1$. The map $(F_{\flat 0}^{\sharp})_{0}$ is calculated similar to (4.40). Because the images of $F_{\flat 0}^{\sharp}$ on the “curvature generators” $\bar{m}^{a}$ and $\bar{\xi}^{\alpha}$ do not belong to
for all $h\in C^{\infty}(M)$ and $s\in\Gamma(\mathfrak{g}[1]^{*})$.
Compatibility with the source maps then reads as
By our usual argument, i.e. Remark A.1, the map $F_{1}^{\prime}$ is given by a lift of a graded map $F_{1\flat}^{\prime}:\mathcal{X}\times\Theta\times\Theta\times\mathcal{N}\to\mathsf{G}$ and (4.50) translates to $\mathsf{t}\circ F_{1\flat}^{\prime}=\mathsf{p}_{M}\circ\psi\circ F_{0}\circ\mathsf{t}$, where $\mathsf{p}_{M}$ is again the projection to the underlying manifold $M$ of $\mathrm{T}[1]M$. Furthermore, as $|\mathscr{P}\mathsf{air}(\Theta)|$ is a trivial Lie groupoid and $|\mathscr{A}^{\mathsf{W}}_{\mathcal{H}}|=(\mathsf{G}\rightrightarrows M)$, the condition $\mathsf{e}\circ F_{0}=F_{1}\circ\mathsf{e}$ implies
$|F_{1}^{\prime}|=|F_{1}|=\mathsf{e}\circ|F_{0}|~{}.$ (4.52)
[…]
and satisfy the identities
[…]
The map $F_{1\flat}^{\prime\prime}$ is now defined by $|F^{\prime\prime}_{1\flat}|$ and a derivation
$H\colon|F^{\prime\prime}_{1\flat}|^{*}C^{\infty}(M)\to C^{\infty}(\mathcal{X})_{1}\otimes C^{\infty}(\mathcal{N})_{0}$ (4.57)
according to
$\begin{gathered}(F^{\prime\prime}_{1\flat})^{\sharp}(g)=|F^{\prime\prime}_{1\flat}|^{*}g-\theta H(|F^{\prime\prime}_{1\flat}|^{*}g)~{}\in~{}C^{\infty}(|\mathcal{X}|\times|\mathcal{N}|)~{}\oplus~{}\theta C^{\infty}(\mathcal{X})_{1}\otimes C^{\infty}(\mathcal{N})_{0}\end{gathered}$ (4.58)
for $g\in C^{\infty}(\mathsf{G})$. As argued in [Li14, Prop. 8.21], any such derivation $H(|F^{\prime\prime}_{1\flat}|^{*}g)$ is given by the de Rham differential followed by a morphism of modules
$H^{\prime}\colon|F^{\prime\prime}_{1\flat}|^{*}\Omega^{1}(\mathsf{G})\to C^{\infty}(\mathcal{X})_{1}\otimes C^{\infty}(\mathcal{N})_{0}~{},$ (4.59)
$H^{\prime}\left(|F^{\prime\prime}_{1\flat}|^{*}\mathsf{t}^{*}\varpi\right)=0$ (4.60)
$H^{\prime\prime}\colon|F^{\prime\prime}_{1\flat}|^{*}\Gamma(\ker(\mathrm{d}\mathsf{t}))^{*}\to C^{\infty}(\mathcal{X})_{1}\otimes C^{\infty}(\mathcal{N})_{0}$ (4.61)
and $\mathsf{i}^{*}$ is induced by $\mathsf{i}\colon\mathsf{Ker}(\mathrm{d}\mathsf{t})\to\mathrm{T}\mathsf{G}$. Finally, since $|F^{\prime\prime}_{\flat 1}|=\mathsf{e}\circ|F|$, we have
$H^{\prime\prime}\colon|F|^{*}\mathsf{e}^{*}\Gamma(\ker(\mathrm{d}\mathsf{t}))^{*}=|F|^{*}\Gamma(\mathfrak{g}[1])^{*}\to C^{\infty}(\mathcal{X})_{1}\otimes C^{\infty}(\mathcal{N})_{0}~{}.$ (4.62)
Therefore, we can identify $H^{\prime\prime}$ with $(\bar{F}_{1})^{\sharp}$.
$\begin{gathered}|F^{\prime}_{1\flat}|:|\mathcal{X}|\times|\mathcal{N}|\to\mathsf{G}~{}~{}~{}\mbox{with}~{}~{}~{}|F^{\prime}_{1\flat}|=\mathsf{e}\circ|F|~{},\\ (F^{\prime}_{1\flat})^{\sharp}(g)=|F^{\prime}_{1\flat}|^{*}(g)-\theta_{2}(\bar{F}_{1})^{\sharp}\circ|F|^{*}(\rho^{*}(\mathsf{e}^{*}\mathrm{d}g))+(\theta_{2}-\theta_{1})(\bar{F}_{1})^{\sharp}\circ|F|^{*}(\mathsf{e}^{*}\mathsf{i}^{*}\mathrm{d}g)\\ -\theta_{2}\theta_{1}\frac{1}{2}[(\bar{F}_{1})^{\sharp}\circ|F|^{*}(\mathsf{e}^{*}\mathsf{i}^{*}\mathrm{d}g),(\bar{F}_{1})^{\sharp}\circ|F|^{*}(\mathsf{e}^{*}\mathsf{i}^{*}\mathrm{d}g)]~{},\end{gathered}$ (4.63)
and the lift of $F_{1\flat}^{\prime}$ to $F_{1}^{\prime}$ is simply
Let us again choose local coordinates $(m_{1}^{a},\bar{m}_{1}^{a},\xi^{\alpha},\bar{\xi}^{\alpha})$ for $\mathrm{T}[1]\mathfrak{g}[1]$ in a neighborhood $U\subset M$. The embedding $\mathsf{e}:M\rightarrow\mathsf{G}$ the allows us to identify $\mathsf{e}(U)\subset\mathsf{G}$ with $U\times V$, on which we use local coordinates $(m_{2}^{a},\bar{m}_{2}^{a},y^{\alpha},\bar{y}^{\alpha})$. In the fibered product $\mathrm{T}[1]\mathfrak{g}[1]\mathbin{{}_{\psi}\times_{\mathrm{d}\mathsf{t}}}\mathrm{T}[1]\mathsf{G}$ we need to identify $\bar{m}_{1}^{a}+\mathtt{r}^{a}_{\alpha}\xi^{\alpha}$ with $\bar{m}_{2}^{a}$ and $m_{1}$ with $m_{2}$. We will show later that
where $O_{2}$ represents functions of order at least 2 in its argument.
$\begin{split}(\bar{F}_{0})^{\sharp}(\xi^{\alpha})+\theta\mathrm{d}_{\mathcal{N}}(\bar{F}^{1})^{\sharp}(\xi^{\alpha})+\theta|F|^{*}(\omega^{\alpha}_{a\beta})F_{0}^{\sharp}(\bar{m}_{1}^{a})(\bar{F}^{1})^{\sharp}(\xi^{\alpha})+\theta|F|^{*}(\mathtt{f}_{\beta\gamma}^{\alpha})(\bar{F}^{1})^{\sharp}(\xi^{\beta})(\bar{F}_{0})^{\sharp}(\xi^{\gamma})~{},\end{split}$ (4.67)
Step 3. Now we prove Equation 4.66. Recall that the Lie algebroid connection corresponding to the Lie groupoid connection is given by [CSS12, Thm. 3.11] (see also Proposition 2.12)
Firstly, we show that
Using Baker–Campbell–Hausdorff formula, one can see that
If one notes that the left hand side is nothing but $\mathrm{d}R_{\mathsf{exp}(-\epsilon\gamma)}\sigma_{\mathcal{H}}\bigl{(}\mathsf{exp}(\epsilon\gamma_{x}),V_{x}\bigr{)}$ with $\epsilon\gamma=y^{\alpha}\bar{y}_{\alpha}$ and $V=\bar{m}$, then it is clear that $\bar{U}(m,y)_{a\beta}^{\alpha}=-\omega_{a\beta}^{\alpha}$.
Recall that in Proposition 2.11 we showed that $\mathrm{Ad}(g^{-1},\gamma)=(g^{-1},0)\circ(\mathsf{e}_{\mathsf{t}(g)},\gamma)\circ\sigma_{\mathcal{H}}(g,\rho(\gamma))$. Using Baker–Campbell–Hausdorff formula for Lie groupoids, one sees the dual of the map $\mathsf{p}_{\mathfrak{g}[1]}\circ\mathrm{Ad}\circ\mathsf{inv}^{\star}:\mathsf{t}^{*}\mathfrak{g}[1]\to\mathfrak{g}[1]$ takes the form
$=\left(m,0,-y+O_{2}(y),0\right)\circ\bigl{(}(m,\bar{m},y,\bar{y})-\sigma_{\mathcal{H}}(m,\bar{m},y,0)\bigr{)}~{},$ (4.75)
then, it is easy to see that this map dualizes to
@@
$\mathbin{\vartriangleleft}^{\sharp}(\xi)(\hat{\nu},\bar{\hat{\nu}};g,\gamma)$ $\coloneqq\xi\big{(}\mathsf{p}_{\mathsf{s}}(\mathrm{Ad}(g^{-1},\hat{\nu})+\vartheta^{\rm tot}(g,\gamma)))~{},$
@@
and as a result $\mathbin{\vartriangleleft}^{\sharp}(\mathrm{d}h)=\mathsf{s}^{*}\mathrm{d}h-\mathbin{\vartriangleleft}^{\sharp}(\rho^{*}\mathrm{d}h)$. Similarly,
@@
We can now regard an adjustment of a Lie groupoid $\mathscr{G}$ as additional data on this groupoid, without referring to its Lie algebroid. First of all, Theorem 4.14 shows that the connection $\nabla$ in the adjustment of the Weil algebra $W_{\nabla}(\mathfrak{g})$ of the corresponding Lie algebroid $\mathfrak{g}$ can be seen as induced by the Cartan connection $\mathcal{H}$ on $\mathscr{G}$. Also, the primitive $\zeta$ of $\nabla$ for covariant adjustments can be seen as originating from a primitive on the groupoid. This follows from the perspective of connections being closed and their curvatures being exact forms in a particular cohomology (a cohomology as developed in [CSS12, Sec. 3.4]), see also [Fis22, Rem. 6.88]. In the latter paper, this fact is discussed in more detail for the special case in which $\mathscr{G}$ is a Lie group bundle, see also Section 7.2.

### v3

@@ 4.3 Adjustment groupoids and Cartan connections
If $\mathcal{H}$ integrates such a covariant adjustment $(\nabla,\zeta)$, then we will write $\mathscr{A}_{(\mathcal{H},\zeta)}^{\mathsf{W}}$ for the corresponding adjustment groupoid.
@@
$\mathbin{\vartriangleleft}^{\sharp}(\xi)(\hat{\nu},\bar{\hat{\nu}};g,\gamma)$ $\coloneqq\xi\big{(}\mathsf{p}_{\mathsf{s}}(\mathrm{Ad}_{\mathcal{H}}(g^{-1},\hat{\nu})+\vartheta^{\rm tot}(g,\gamma)))~{},$
@@
and as a result, $\mathbin{\vartriangleleft}^{\sharp}(\mathrm{d}h)=\mathsf{s}^{*}\mathrm{d}h-\mathbin{\vartriangleleft}^{\sharp}(\rho^{*}\mathrm{d}h)$. Similarly,
@@
Note that Cartan connections on Lie groupoids can be linearized to Cartan connections on Lie algebroids. Also, the primitive $\zeta$ of $\nabla$ for covariant adjustments can be seen as originating from a primitive on the groupoid. This follows from the perspective of connections being closed and their curvatures being exact forms in a particular cohomology (a cohomology as developed in [CSS12, Sec. 3.4]), see also [Fis21b] and [Fis22, Rem. 6.88]. In the latter paper, this fact is discussed in more detail for the special case in which $\mathscr{G}$ is a Lie group bundle, see also Section 7.2. This shows how the global theory leads to the local theory.
Proposition 4.13.
An adjustment groupoid $\mathscr{A}^{\mathsf{W}}$ of a Lie groupoid $\mathscr{G}$ induces a corresponding Cartan connection on $\mathscr{G}$.
Proof.
Consider the dg-Lie groupoid $\mathscr{A}^{\mathsf{W}}$ satisfying the axioms of Definition 4.7. Using the fact that $\mathsf{t}\colon\mathscr{A}^{\mathsf{W}}_{1}\to\mathrm{T}[1]\mathfrak{g}[1]$ is a surjective submersion [^41] This implies that it has local sections. and the diagram (4.14), one can show that $\mathscr{A}^{\mathsf{W}}_{1}$ is locally isomorphic to $\mathrm{T}[1]\mathfrak{g}[1]\mathbin{{}_{\psi}\times_{\mathrm{d}\mathsf{t}}}\mathrm{T}[1]\mathsf{G}$. On the other hand,
$\big{|}\mathscr{A}^{\mathsf{W}}_{1}\big{|}=\big{|}\mathrm{T}[1]\mathfrak{g}[1]\mathbin{{}_{\psi}\times_{\mathrm{d}\mathsf{t}}}\mathrm{T}[1]\mathsf{G}\big{|}=\mathsf{G}~{},$ (4.43)
and altogether $\mathscr{A}^{\mathsf{W}}_{1}$ and $\mathrm{T}[1]\mathfrak{g}[1]\mathbin{{}_{\psi}\times_{\mathrm{d}\mathsf{t}}}\mathrm{T}[1]\mathsf{G}$ are isomorphic.
Consider the source map $\mathsf{s}\colon\mathrm{T}[1]\mathfrak{g}[1]\mathbin{{}_{\psi}\times_{\mathrm{d}\mathsf{t}}}\mathrm{T}[1]\mathsf{G}\to\mathrm{T}[1]\mathfrak{g}[1]$. As this map is a dg-map, it is given by a graded map $\mathsf{s}_{\flat}\colon\mathrm{T}[1]\mathfrak{g}[1]\mathbin{{}_{\psi}\times_{\mathrm{d}\mathsf{t}}}\mathrm{T}[1]\mathsf{G}\to\mathfrak{g}[1]$. Furthermore, the body $|\mathsf{s}|$ of the map $\mathsf{s}$ must coincide with the source function $\mathsf{s}\colon\mathsf{G}\to M$.
Recall that the sheaf of functions $\mathcal{O}_{\mathfrak{g}[1]}$ is generated by the space $\Gamma(\mathfrak{g}[1]^{*})$. Therefore, the map of algebras $\mathsf{s}^{\sharp}\colon\mathcal{O}_{\mathfrak{g}[1]}\to\mathcal{O}_{\mathrm{T}[1]\mathfrak{g}[1]\mathbin{{}_{\psi}\times_{\mathrm{d}\mathsf{t}}}\mathrm{T}[1]\mathsf{G}}$ is given by the map of modules $\mathsf{s}^{\sharp}_{1}\colon\Gamma(\mathfrak{g}[1]^{*})\to\left(\mathcal{O}_{\mathrm{T}[1]\mathfrak{g}[1]\mathbin{{}_{\psi}\times_{\mathrm{d}\mathsf{t}}}\mathrm{T}[1]\mathsf{G}}\right)_{1}$. Moreover, one can show that
$\left(\mathcal{O}_{\mathrm{T}[1]\mathfrak{g}[1]\mathbin{{}_{\psi}\times_{\mathrm{d}\mathsf{t}}}\mathrm{T}[1]\mathsf{G}}\right)_{1}\cong\left(\mathcal{O}_{(\mathrm{T}[1]M\oplus\mathfrak{g}[1])\times_{\mathrm{T}[1]M}\mathrm{T}[1]\mathsf{G}}\right)_{1}\cong\left(\mathcal{O}_{\mathsf{t}^{*}\mathfrak{g}[1]\oplus\mathrm{T}[1]\mathsf{G}}\right)_{1}~{},$ (4.44)
where $(-)_{1}$ denotes the truncation to the degree 1 elements. We conclude that the map $\mathsf{s}^{\sharp}_{1}$ is given by a map of vector bundles $\left(\mathsf{s}^{\sharp}_{1}\right)^{\star}\colon\mathsf{t}^{*}\mathfrak{g}\oplus\mathrm{T}\mathsf{G}\to\mathfrak{g}$.
The functor $\Psi$ in Definition 4.7 shows that the diagram
[…]
is commutative.
Define the vector bundle map $\omega\colon\mathrm{T}\mathsf{G}\to\mathsf{t}^{*}\mathfrak{g}\oplus\mathrm{T}\mathsf{G}\to\mathfrak{g}$, where the first map is the obvious inclusion and the second map is $\left(\mathsf{s}^{\sharp}_{1}\right)^{\star}$. The diagram above shows that $\omega^{*}\colon\mathrm{T}\mathsf{G}\to\mathsf{s}^{*}\mathfrak{g}$ defines a splitting of the first exact sequence in (2.16). In other words, $\omega^{*}$ is a connection 1-form and the projection $\mathscr{L}\circ\omega^{*}$ to $\ker(\mathrm{d}\mathsf{t})$ defines a distribution $\mathscr{H}$. Note that the map $\omega^{*}$ as splitting of the exact sequence is equivalent to the horizontal lift $\sigma_{\mathcal{H}}\colon\mathsf{t}^{*}\mathrm{T}M\to\mathrm{T}\mathsf{G}$, and we have $\omega^{*}\circ\sigma_{\mathcal{H}}=0$.
Define the map $\mathrm{Ad}^{-1}\colon\mathsf{t}^{*}\mathfrak{g}\to\mathsf{t}^{*}\mathfrak{g}\oplus\mathrm{T}\mathsf{G}\to\mathfrak{g}$, where the first map is the obvious inclusion and the second one is $\left(\mathsf{s}^{\sharp}_{1}\right)^{\star}$. By linearity of the map $\left(\mathsf{s}^{\sharp}_{1}\right)^{\star}$ and the diagram above, we have
$\begin{split}\mathbin{\blacktriangleleft}(\mathsf{t}(g),\nu;\sigma_{\mathcal{H}}(g;\mathsf{t}(g),\rho(\nu)))&=\mathrm{Ad}^{-1}(\mathsf{t}(g),\nu;g)+\omega^{*}\bigl{(}\sigma_{\mathcal{H}}(g;\mathsf{t}(g),\rho(\nu))\bigr{)}\\ &=\mathrm{Ad}^{-1}(\mathsf{t}(g),\nu;g)\\ &=\mathrm{d}L_{g^{-1}}(\mathsf{e}_{\mathsf{t}(g)},\nu)\circ\sigma_{\mathcal{H}}(g;\mathsf{t}(g),\rho(\nu))\\ &=\mathrm{Ad}_{\mathcal{H}}\left(g^{-1};\mathsf{t}(g),\nu\right)~{},\end{split}$ (4.46)
where we have used Proposition 2.10 and the identity $\omega^{*}\circ\sigma_{\mathcal{H}}=0$, and $\mathrm{Ad}_{\mathcal{H}}$ is the adjoint action defined in Definition 2.9.
By linearity of $\left(\mathsf{s}^{\sharp}_{1}\right)^{\star}$, it follows that this map coincides with the one in Equation 4.19. The connection $\mathcal{H}$ has to be Cartan as $\mathscr{A}^{\mathsf{W}}$ forms a groupoid. ∎